OpenStudy (anonymous):

@zepdrix 0.007(3-k)+0.04(k-8)=1

zepdrix (zepdrix):

I guess we might as well do the same thing we did last time. Get rid of the decimals (or fractions) first. This is a little trickier. Let's convert our decimal to a fraction. It looks like our first one is in the thousandth place. 7 thousandths. While the 4 is in the hundredth place. 4 hundredths.\[\large \frac{7}{1000}(3-k)+\frac{4}{100}(k-8)=1\] So our largest denominator is 1000, let's multiply both sides by 1000, giving us,\[\large 7(3-k)+40(k-8)=1000\] Understand how I changed the decimals? that part is a bit tough.

OpenStudy (anonymous):

yep!

zepdrix (zepdrix):

So from here it's not too bad. Distribute the 7 to each term in the first brackets,\[\large \color{orangered}{21-7k}+40(k-8)=1000\]Distribute the 40 to each term in the second set of brackets,\[\large 21-7k+\color{orangered}{40k-320}=1000\]Combine like terms,\[\large 33k-299=1000\]Adding 299 to each side gives us,\[\large 33k=1299\]Diving both sides by 33 gives us,\[\large k=\frac{1299}{33}\] Which is approximately, \(\large 39.36\).

OpenStudy (anonymous):

well, i had to reduece it to 433/11 and then it was right in math lab (im going to try the next one on my own; ill contact you if i cant get it